A solution containing 80.0 grams of glucose (C6H12O6) is dissolved in enough water to make 250 mL of solution. What is its molarity?


1
Expert's answer
2021-08-04T07:13:24-0400

Molar mass of glucose = 180.156 


Moles of glucose "=\frac{80}{180.156}=0.444moles"


Volumes = "\frac{250}{1000}=0.25"


"Molarity=\frac{0.25}{0.444}=0.56M"